A window is made up of a square portion and an equilateral triangle portion above it. The base of
the triangular portion coincides with the upper side of the square. If the perimeter of the window is
6 m, the area of the window in m2
is ___________.

Answer by rfer(16322)

Triangle=
a=1/2LW
a=0.5*1.2*1.2
a=0.72 m2
--------------
Square
a=LW
a=1.2*1.2
a=1.44 m2
-----------
total
a=1.44+0.72=2.16 m2